# Theoretical Detailed Analyses for DC readout and a Fabri-Pérot gravitational-wave detector

### v1

@@ Appendix A Power counting photodetection
Through the original definition (LABEL:eq:power-operator-quadrature-Fourier) of the power operator is different from the definition of Glauber’s photon number (III), in the situation $\omega_{0}\gg\Omega$, the tedious but straightforward calculations lead to the expression of the stationary noise spectral density including the one-point support function defined by Eq. (33). As in the case of Glauber’s photon number, using the formulae
@@
we reach the expression
@@
where we used the sideband picture (44) and the situation $\Omega\ll\omega_{0}$. Furthermore, we introduce the amplitude- and the phase-quadratures $\hat{b}_{n1}(\Omega)$ and $\hat{b}_{n2}(\Omega)$ by Eqs. (45) and (46). Moreover, we define the operator $\hat{b}_{n\Theta}(\Omega)$ defined by Eq. (47). Then, we obtain
@@ Appendix B Explicit evaluation of the radiation pressure forces to the mirrors
In this appendix, we show the evaluation of the radiation pressure forces on the mirrors in Sec. V.2. The crucial premise of this evaluation is that the radiation pressure forces on the mirrors are determined by the power of the laser, which corresponds to the pointing flux of the laser that affects the mirrors. As described in Ref. [8], the power operator of the laser is given by ${\cal A}\hat{E}^{2}(t)/(4\pi)$, where $\hat{E}(t)$ is the electric field operator which touch to mirrors and $\cal A$} is the cross-sectional area of the optical beam which introduced in Eq. (3).
@@
We also note that we have to evaluate the Fourier transformations of the mirror displacement operators $D_{d}^{\dagger}\hat{Z}_{x}D_{d}$, $D_{d}^{\dagger}\hat{Z}_{y}D_{d}$, $D_{d}^{\dagger}\hat{Z}_{XITM}D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{YITM}D_{d}$ instead of operators $\hat{Z}_{x}$, $\hat{Z}_{y}$, $\hat{Z}_{XITM}$, and $\hat{Z}_{YITM}$ when we evaluate the input-output relation (117). Here, the operator $D_{d}$ is the displacement operator associated with the coherent state for the optical quadrature $\hat{D}(\omega):=\hat{d}(\omega)\Theta(\omega)+\hat{d}^{\dagger}(-\omega)\Theta(-\omega)$ which is defined by Eq. (101). For the quadrature $\hat{D}(\omega)$, $D_{d}^{\dagger}\hat{D}(\omega)D_{d}$ are given by Eq. (104). Trivially, the quadrature $\hat{A}(\omega)$ commutes with the displacement operator $D_{d}$. Keep in our mind the properties of the displacement operator $D_{d}$, we evaluate the radiation pressure forces (326), (327), (329), and (331) by the operations $D_{d}^{\dagger}$ from the left and $D_{d}$ from the right. Furthermore, we ignore the terms include $\hat{D}_{v}(\omega)\hat{D}_{v}(\omega^{\prime})$, $\hat{D}_{v}(\omega)\hat{A}(\omega^{\prime})$, and $\hat{A}(\omega)\hat{A}(\omega^{\prime})$. We symbolically denote these terms $\hat{D}_{v}^{2}$, $\hat{D}_{v}\hat{A}$, and $\hat{A}^{2}$. Moreover, we also ignored the terms that include $\hat{D}_{v}(\omega)\hat{Z}_{x}(\omega^{\prime})$, $\hat{D}_{v}(\omega)\hat{Z}_{XITM}(\omega^{\prime})$, $\hat{D}_{v}(\omega)\hat{Z}_{y}(\omega^{\prime})$, $\hat{D}_{v}(\omega)\hat{Z}_{YITM}(\omega^{\prime})$, $\hat{A}(\omega)\hat{Z}_{x}(\omega^{\prime})$, $\hat{A}(\omega)\hat{Z}_{XITM}(\omega^{\prime})$, $\hat{A}(\omega)\hat{Z}_{y}(\omega^{\prime})$, $\hat{A}(\omega)\hat{Z}_{YITM}(\omega^{\prime})$. We symbolically denote these terms as $\hat{D}_{v}\hat{X}$ and $\hat{A}\hat{X}$. Through these evaluations, the radiation pressure forces (326), (327), (329), and (331) are given by
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\omega t}D_{d}^{\dagger}\hat{F}_{rpXEM}(t)D_{d}$ (332)
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{XITM}(\omega-\omega_{1}-\omega_{2})D_{d}$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{x}(\omega-\omega_{1}-\omega_{2})D_{d}$
@@
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\omega t}D_{d}^{\dagger}\hat{F}_{rpYEM}(t)D_{d}$ (333)
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{YITM}(\omega-\omega_{1}-\omega_{2})D_{d}$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{y}(\omega-\omega_{1}-\omega_{2})D_{d}$
@@
$\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}D_{d}^{\dagger}\hat{F}_{rpXITM}(t)D_{d}$ (334)
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{XITM}(\omega-\omega_{1}-\omega_{2})D_{d}$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{x}(\omega-\omega_{1}-\omega_{2})D_{d}$
@@
$\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}D_{d}^{\dagger}\hat{F}_{rpYITM}(t)D_{d}$ (335)
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{YITM}(\omega-\omega_{1}-\omega_{2})D_{d}$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2})D_{d}^{\dagger}\hat{Z}_{y}(\omega-\omega_{1}-\omega_{2})D_{d}$
@@
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\Omega t}D_{d}^{\dagger}\hat{F}_{rpXEM}(t)D_{d}$ (336)
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{XITM}(\Omega)D_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{x}(\Omega)D_{d}$
@@
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\Omega t}D_{d}^{\dagger}\hat{F}_{rpYEM}(t)D_{d}$ (337)
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{YITM}(\Omega)D_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{y}(\Omega)D_{d}$
@@
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega)D_{d}$ (338)
$=$ $\int_{-\infty}^{+\infty}dte^{+i\Omega t}\frac{1}{m_{ITM}}D_{d}^{\dagger}\hat{F}_{rpXITM}(t)D_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]D_{d}^{\dagger}\hat{Z}_{XITM}(\Omega)D_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{x}(\Omega)D_{d}$
@@
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega)D_{d}$ (339)
$=$ $\int_{-\infty}^{+\infty}dte^{+i\Omega t}\frac{1}{m_{ITM}}D_{d}^{\dagger}\hat{F}_{rpYITM}(t)D_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]D_{d}^{\dagger}\hat{Z}_{YITM}(\Omega)D_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{y}(\Omega)D_{d}$
@@
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega)D_{d}$ (340)
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]D_{d}^{\dagger}\hat{Z}_{XITM}(\Omega)D_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{x}(\Omega)D_{d}$
@@
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega)D_{d}$ (341)
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]\left[\frac{1}{m_{EM}}e^{+i\Omega\tau}+\frac{1}{m_{ITM}}\right]D_{d}^{\dagger}\hat{Z}_{YITM}(\Omega)D_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}D_{d}^{\dagger}\hat{Z}_{y}(\Omega)D_{d}$
@@
In these expressions (338), (339), (340), and (341), we have the terms proportional to $D_{d}^{\dagger}\hat{Z}_{XITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{x}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{YITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z_{y}}(\Omega)D_{d}$ which includes the factor $\sin(2\omega_{0}\tau)$. These terms correspond to the optical spring effects of the Fabri-Pérot interferometer [32, 33]. However, these optical spring effects can be ignored through the tuning condition (161) due to the factor $\sin(2\omega_{0}\tau)$. In this paper, we only consider the situation of this tuning, in which there is no optical spring effect.

### v2

@@ Appendix A Power counting photodetection
Through the original definition (LABEL:eq:power-operator-quadrature-Fourier) of the power operator is different from the definition of Glauber’s photon number (III), in the situation $\omega_{0}\gg\Omega$, the tedious but straightforward calculations lead to the expression of the stationary noise spectral density including the one-point support function defined by Eq. (33). As in the case of Glauber’s photon number, through the final input-output relation (189) which was derived in Sec. VI, we obtain the formulae
@@
in the same way to obtain Eqs. (39)–(42) in Sec. II.2. Using these formulae, (318)–(321) we reach the expression
@@
where we used the sideband picture (44) and the situation $\Omega\ll\omega_{0}$.
Furthermore, we introduce the amplitude- and the phase-quadratures $\hat{b}_{n1}(\Omega)$ and $\hat{b}_{n2}(\Omega)$ by Eqs. (45) and (46). Moreover, we define the operator $\hat{b}_{n\Theta}(\Omega)$ defined by Eq. (47). Then, we obtain
@@ Appendix B Explicit evaluation of the radiation pressure forces to the mirrors
In this appendix, we show the evaluation of the radiation pressure forces on the mirrors in Sec. V.2. The crucial premise of this evaluation is that the radiation pressure forces on the mirrors are determined by the laser’s power, which corresponds to the laser’s pointing flux and affects the mirrors. As described in Ref. [10], the power operator of the laser is given by ${\cal A}\hat{E}^{2}(t)/(4\pi)$, where $\hat{E}(t)$ is the electric field operator which touch to mirrors and $\cal A$} is the cross-sectional area of the optical beam which introduced in Eq. (3).
@@
We also note that we have to evaluate the Fourier transformations of the mirror displacement operators ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}{{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{y}{{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}{{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}{{\mathfrak{D}}}_{d}$ instead of operators $\hat{Z}_{x}$, $\hat{Z}_{y}$, $\hat{Z}_{XITM}$, and $\hat{Z}_{YITM}$ when we evaluate the input-output relation (117). Here, the operator ${{\mathfrak{D}}}_{d}$ is the displacement operator associated with the coherent state for the optical quadrature $\hat{D}(\omega):=\hat{d}(\omega)\Theta(\omega)+\hat{d}^{\dagger}(-\omega)\Theta(-\omega)$ which is defined by Eq. (101). For the quadrature $\hat{D}(\omega)$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{D}(\omega){{\mathfrak{D}}}_{d}$ are given by Eq. (104). Trivially, the quadrature $\hat{A}(\omega)$ commutes with the displacement operator ${{\mathfrak{D}}}_{d}$. Keep in our mind the properties of the displacement operator ${{\mathfrak{D}}}_{d}$, we evaluate the radiation pressure forces (326), (327), (329), and (331) by the operations ${{\mathfrak{D}}}_{d}^{\dagger}$ from the left and ${{\mathfrak{D}}}_{d}$ from the right. Furthermore, we ignore the terms include $\hat{D}_{v}(\omega)\hat{D}_{v}(\omega^{\prime})$, $\hat{D}_{v}(\omega)\hat{A}(\omega^{\prime})$, and $\hat{A}(\omega)\hat{A}(\omega^{\prime})$. We symbolically denote these terms $\hat{D}_{v}^{2}$, $\hat{D}_{v}\hat{A}$, and $\hat{A}^{2}$. Moreover, we also ignored the terms that include $\hat{D}_{v}(\omega)\hat{Z}_{x}(\omega^{\prime})$, $\hat{D}_{v}(\omega)\hat{Z}_{XITM}(\omega^{\prime})$, $\hat{D}_{v}(\omega)\hat{Z}_{y}(\omega^{\prime})$, $\hat{D}_{v}(\omega)\hat{Z}_{YITM}(\omega^{\prime})$, $\hat{A}(\omega)\hat{Z}_{x}(\omega^{\prime})$, $\hat{A}(\omega)\hat{Z}_{XITM}(\omega^{\prime})$, $\hat{A}(\omega)\hat{Z}_{y}(\omega^{\prime})$, $\hat{A}(\omega)\hat{Z}_{YITM}(\omega^{\prime})$. We symbolically denote these terms as $\hat{D}_{v}\hat{X}$ and $\hat{A}\hat{X}$. Through these evaluations, the radiation pressure forces (326), (327), (329), and (331) are given by
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\omega t}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpXEM}(t){{\mathfrak{D}}}_{d}$ (332)
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
@@
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\omega t}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpYEM}(t){{\mathfrak{D}}}_{d}$ (333)
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\omega-\omega_{1})\tau}\right]^{-1}\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{y}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
@@
$\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpXITM}(t){{\mathfrak{D}}}_{d}$ (334)
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
@@
$\int_{-\infty}^{+\infty}dte^{+i\omega t}\frac{1}{m_{ITM}}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpYITM}(t){{\mathfrak{D}}}_{d}$ (335)
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\hat{D}_{c}(\omega_{1})\hat{D}_{c}(\omega_{2}){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{y}(\omega-\omega_{1}-\omega_{2}){{\mathfrak{D}}}_{d}$
@@
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\Omega t}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpXEM}(t){{\mathfrak{D}}}_{d}$ (336)
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}(\Omega){{\mathfrak{D}}}_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}(\Omega){{\mathfrak{D}}}_{d}$
@@
$\frac{1}{m_{EM}}\int_{-\infty}^{+\infty}dte^{+i\Omega t}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpYEM}(t){{\mathfrak{D}}}_{d}$ (337)
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}(\Omega){{\mathfrak{D}}}_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{y}(\Omega){{\mathfrak{D}}}_{d}$
@@
${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega){{\mathfrak{D}}}_{d}$ (338)
$=$ $\int_{-\infty}^{+\infty}dte^{+i\Omega t}\frac{1}{m_{ITM}}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpXITM}(t){{\mathfrak{D}}}_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}(\Omega){{\mathfrak{D}}}_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}(\Omega){{\mathfrak{D}}}_{d}$
@@
${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega){{\mathfrak{D}}}_{d}$ (339)
$=$ $\int_{-\infty}^{+\infty}dte^{+i\Omega t}\frac{1}{m_{ITM}}{{\mathfrak{D}}}_{d}^{\dagger}\hat{F}_{rpYITM}(t){{\mathfrak{D}}}_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}(\Omega){{\mathfrak{D}}}_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{y}(\Omega){{\mathfrak{D}}}_{d}$
@@
${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega){{\mathfrak{D}}}_{d}$ (340)
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}(\Omega){{\mathfrak{D}}}_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}(\Omega){{\mathfrak{D}}}_{d}$
@@
${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega){{\mathfrak{D}}}_{d}$ (341)
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-e^{+2i\Omega\tau}\right]\left[\frac{1}{m_{EM}}e^{+i\Omega\tau}+\frac{1}{m_{ITM}}\right]{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}(\Omega){{\mathfrak{D}}}_{d}$
@@
$\quad\quad\times\left[1-\sqrt{1-T}e^{+2i(\Omega+\omega_{0})\tau}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i(\Omega-\omega_{0})\tau}\right]^{-1}{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{y}(\Omega){{\mathfrak{D}}}_{d}$
@@
In these expressions (338), (339), (340), and (341), we have the terms proportional to ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{XITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{x}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{YITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z_{y}}(\Omega){{\mathfrak{D}}}_{d}$ which includes the factor $\sin(2\omega_{0}\tau)$. These terms correspond to the optical spring effects of the Fabry-Pérot interferometer [37, 38]. However, these optical spring effects can be ignored through the tuning condition (161) due to the factor $\sin(2\omega_{0}\tau)$. In this paper, we only consider the situation of this tuning, in which there is no optical spring effect.